Obtain the Fourier series for the function f whose definition in one period is f(x)= -x, -1≤x≤1.Sketch the graph of f
1
Expert's answer
2020-05-18T19:06:57-0400

"f(x)=-x, x\in[-1,1]"

"f(x)=\frac{a_0}{2}+\sum\limits^{\infty}_{m=1}(a_m\cos(m\pi x)+b_m\sin(m\pi x))"

where

"a_m=\int\limits^{1}_{-1}(-x)\cos(m\pi x)dx, m=0,1,2,...\\
b_m=\int\limits^{1}_{-1}(-x)\sin(m\pi x)dx, m=1,2,..."

Let "m=0"

"a_0=\int\limits^{1}_{-1}(-x)dx=-\frac{x^2}{2}|^{1}_{-1}=-\frac{1}{2}+\frac{1}{2}=0\\
a_m=\int\limits^{1}_{-1}(-x)\cos(m\pi x)dx=-\frac{x}{\pi m}\sin( m\pi x)|^{1}_{-1}+\\
+\frac{1}{\pi m}\int\limits^{1}_{-1}\sin(m\pi x)dx=-\frac{1}{(\pi m)^2}\cos (m\pi x)|^{1}_{-1}=\\
=-\frac{1}{(\pi m)^2}(\cos m\pi- \cos m\pi)=0\\
b_m=\int\limits^{1}_{-1}(-x)\sin(m\pi x)dx=\frac{x}{\pi m}\cos (m\pi x)|^{1}_{-1}-\\
-\frac{1}{\pi m}\int\limits^{1}_{-1}\cos(m\pi x)dx=\frac{2}{\pi m}\cos m\pi -\\
-\frac{1}{(\pi m)^2}\sin (m\pi x)|^{1}_{-1}=\frac{2}{\pi m}(-1)^m\\"


"-x=\sum\limits^{\infty}_{m=1}\frac{2}{\pi m}(-1)^m\sin(m\pi x)\\"